An FI has purchased (borrowed) a one-year $10 million Eurodollar deposit at an annual interest rate of 6 percent. It has invested these proceeds in one-year Euro (€) bonds at an annual rate of 6.5 percent after converting them at the current spot rate of €1.75/$. Both interest and principal are paid at the end of the year.

What is the spread earned by the bank at the end of the year if the exchange rate remains at €1.75/$? 


A. 0.50 percent.


B. 1.00 percent.


C. 1.5 percent.


D. 2.0 percent.


E. 2.5 percent.

An FI has purchased (borrowed) a one-year $10 million Eurodollar deposit at an annual interest rate of 6 percent. It has invested these proceeds in one-year Euro (€) bonds at an annual rate of 6.5 percent after converting them at the current spot rate of €1.75/$. Both interest and principal are paid at the end of the year.

What is the spread earned by the bank if the end-of-year exchange rate is €1.77/$?
A. -1.00 percent.
B. -0.70 percent.
C. -0.25 percent.
D. 0.00 percent.
E. 0.20 percent.

103. An FI has purchased (borrowed) a one-year $10 million Eurodollar deposit at an annual interest rate of 6 percent. It has invested these proceeds in one-year Euro (€) bonds at an annual rate of 6.5 percent after converting them at the current spot rate of €1.75/$. Both interest and principal are paid at the end of the year.

What is the spread earned if the bank can sell one-year forward Euros at €1.755/$?
A. -0.70 percent.
B. -0.25 percent.
C. 0.00 percent.
D. 0.20 percent.
E. 0.50 percent.

An FI has purchased (borrowed) a one-year $10 million Eurodollar deposit at an annual interest rate of 6 percent. It has invested these proceeds in one-year Euro (€) bonds at an annual rate of 6.5 percent after converting them at the current spot rate of €1.75/$. Both interest and principal are paid at the end of the year.

Assume that instead of investing in Euro bonds at a fixed rate of 6.5 percent, it invests them in variable rates of LIBOR + 1.5 percent, reset every six months. The current LIBOR rate is 5 percent. What is the annual spread earned by the bank if LIBOR at the end of six months is 5.5 percent? Assume both interest and principal will be reinvested in six months. Assume the exchange rate remains at €1.75/$ at the end of the year.
A. 0.50 percent.
B. 0.68 percent.
C. 0.86 percent.
D. 0.90 percent.
E. 0.95 percent.

An FI has purchased (borrowed) a one-year $10 million Eurodollar deposit at an annual interest rate of 6 percent. It has invested these proceeds in one-year Euro (€) bonds at an annual rate of 6.5 percent after converting them at the current spot rate of €1.75/$. Both interest and principal are paid at the end of the year.

Assume that instead of investing in Euro bonds at a fixed rate of 6.5 percent, the FI invests them in variable rates of LIBOR + 1.5 percent, reset every six months. The current LIBOR rate is 5 percent. Assume both interest and principal will be reinvested in six months. Assume the exchange rate remains at €1.75/$ at the end of the year. What should be the LIBOR rates in six months in order for the bank to earn a 1 percent spread?
A. 5.25 percent.
B. 5.48 percent.
C. 5.76 percent.
D. 5.86 percent.
E. 5.94 percent.
Answer #1

1). Amount borrowed ($) = 10 mn; Amount to be returned ($) = 10*(1+6%) = 10.6 mn

Amount invested (Euro) = E1.75/$*10 = 17.5 mn; Amount earned (Euro) = 17.5*(1+6.5%) = 18.64 mn

Amount earned ($) = 18.64/E1.75/$ = 10.65 mn

Spread ($) = Amount earned ($) - Amount to be returned ($) = 10.65 - 10.6 = 0.05 mn

%age spread = spread ($)/amount borrowed ($) = 0.05/10 = 0.50% (option A)

2). Amount borrowed ($) = 10 mn; Amount to be returned ($) = 10*(1+6%) = 10.6 mn

Amount invested (Euro) = E1.75/$*10 = 17.5 mn; Amount earned (Euro) = 17.5*(1+6.5%) = 18.64 mn

Amount earned ($) = 18.64/E1.77/$ = 10.53 mn

Spread ($) = Amount earned ($) - Amount to be returned ($) = 10.53 - 10.6 = -0.07

%age spread = spread ($)/amount borrowed ($) = -0.07/10 = -0.70% (option B)

3). Amount borrowed ($) = 10 mn; Amount to be returned ($) = 10*(1+6%) = 10.6 mn

Amount invested (Euro) = E1.75/$*10 = 17.5 mn; Amount earned (Euro) = 17.5*(1+6.5%) = 18.64 mn

Amount earned ($) = 18.64/E1.76/$ = 10.62 mn

Spread ($) = Amount earned ($) - Amount to be returned ($) = 10.62 - 10.6 = 0.02 mn

%age spread = spread ($)/amount borrowed ($) = 0.02/10 = 0.20% (option D)

4). Amount borrowed ($) = 10 mn; Amount to be returned ($) = 10*(1+6%) = 10.6 mn

Amount invested (Euro) = E1.75/$*10 = 17.5 mn; Amount earned (Euro) = [17.5*(1+6.5%/2)]*(1+7%/2) = 18.70 mn

Amount earned ($) = 18.70/E1.75/$ = 10.69 mn

Spread ($) = Amount earned ($) - Amount to be returned ($) = 10.69 - 10.6 = 0.09 mn

%age spread = spread ($)/amount borrowed ($) = 0.09/10 = 0.86% (option A)

5). %age spread = 1% so spread ($) = 1%*amount invested = 1%*10 = 0.10 mn

Amount borrowed ($) = 10 mn; Amount to be returned ($) = 10*(1+6%) = 10.6 mn

Amount to be earned ($) = Amount borrowed + spread = 10.6 + 0.1 = 10.7 mn

Amount to be earned (Euro) = Euro1.75/$*10.7 = 18.725 mn

Amount invested (Euro) = 17.5 mn so amount after 6 months (Euro) = 17.5*(1+6.5%/2) = 18.07 mn

Let total rate for the next six months be r%. Then

18.07*(1+ r%/2) = 18.725 mn

1+r%/2 = (18.725/18.07) = 1.0363

r% = 7.26% so LIBOR is 7.26% - 1.5% = 5.76% (option C)

Note: There may be minor differences in numbers due to rounding off.